a new compound was recently discovered and found to have an automatic weight of 342.38 amu


1
Expert's answer
2021-10-24T01:28:27-0400

average atomic mass = ( % age abundance of lighter isotope × its atomic mass) + (% age abundance of heavier isotope × its atomic mass) / 100


percentage of heavier isotope = 100- 68.322 = 31.678


Now we will put the values in formula.


342.38 = (68.322× 340.91) + (31.678 × X) / 100


342.38 = 23291.65302 + (31.678 × X) / 100


342.38 × 100 = 23291.65302 + (31.678 × X)


34238 -23291.65302 = (31.678 × X)


10946.35 / 31.678 = X


345.6 = X


The mass of heavier isotope is 345.6 amu.